slipknot3066:

Probability! I'm moving up in the realm of Mathematics. Anita and Amy appeared for the same interview. Depending on their qualifications, the probability that Anita will be selected is 5/7 and the probability that Amy will be selected is 2/5

slipknot3066:

1) The probability that neither of them is selected is:_____________ and ________________ 2) The probability that at least one of them is selected is: _____________ and ______________

slipknot3066:

@Narad @Angle If you two can help just a lil bit I'd be in you're debt.

Narad:

The probability that Anita is selected is = p =5/7 The probability that Amy is selected is =q= 2/5 The probability that neither of them is selected is =(1-p) or (1-q) The probability that at least one of them is selected is either Anita is selected and Amy is not selected = p(1-q) or Amy is selected and Anita is not selected =q(1-p)

dude:

the probability that neither of them is accepted if P(Anita not selected)*P(amy nont selected) if that helps The probability of either one not being selected is 1-P(selected) for the second bit, the probability that at least one is selected can be thought of as P(amy selected and anita is not)+P(anita slected and amy is not)+P(both selected) Credits to @kittybasil

kittybasil:

Thanks for that :) (the original credit is someone named Tommo off a math chat lol)